OpenStudy (anonymous):

Beacon Electronics advertised a portable TV set on sale for $272.49 after a 35% markdown from the original price. What was the original price of the TV set? (Round to the nearest cent.)

OpenStudy (anonymous):

272.49 = 0.65 P Solve for P

OpenStudy (anonymous):

i did i got it wrong

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

the problem i need the correct ansewr

OpenStudy (jdoe0001):

$272.49 is the price after the markdown of 35% so $272.49 is really the 65% ( 100% - 35% = 65%) so say the original amount was really "x", thus \(\large \begin{array}{cclll} quantity&percentage\\ \hline\\ 272.49&65\\ x&100 \end{array}\implies \cfrac{272.49}{x}=\cfrac{65}{100}\)

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

is that right

OpenStudy (jdoe0001):

well... think about it if the TV on sale is $272.49 and is discounted by 35% already the original price would be more than that though it'd be nice to find a TV for $1.50

OpenStudy (anonymous):

lol so what is it

OpenStudy (jdoe0001):

solve for "x", cross-multiply

OpenStudy (anonymous):

idk im get confuse

OpenStudy (jdoe0001):

\(\large{\begin{array}{cclll} quantity&percentage\\ \hline\\ 272.49&65\\ x&100 \end{array}\implies} \cfrac{272.49}{{\color{blue}{ x}}}=\cfrac{65}{{\color{blue}{ 100}}} \\ \quad \\ cross-multiply\implies 272.49\cdot {\color{blue}{ 100}}=\cdot 65\implies \cfrac{272.49\cdot {\color{blue}{ 100}}}{65}={\color{blue}{ x}}\)

OpenStudy (jdoe0001):

shoot.. I have a typo.. one sec

OpenStudy (anonymous):

k

OpenStudy (jdoe0001):

\(\large{\begin{array}{cclll} quantity&percentage\\ \hline\\ 272.49&65\\ x&100 \end{array}\implies} \cfrac{272.49}{{\color{blue}{ x}}}=\cfrac{65}{{\color{blue}{ 100}}} \\ \quad \\ cross-multiply\implies 272.49\cdot {\color{blue}{ 100}}={\color{blue}{ x}}\cdot 65\implies \cfrac{272.49\cdot {\color{blue}{ 100}}}{65}={\color{blue}{ x}}\)

OpenStudy (anonymous):

419.21

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

it says round to nearest cent

OpenStudy (jdoe0001):

419.21538461538461538462 <---- round to the nearest 100th, thus cents

OpenStudy (jdoe0001):

notice \(\bf 419.21{\color{blue}{ 5}}38461538461538462\)

OpenStudy (anonymous):

419

OpenStudy (jdoe0001):

well, that'd be rounded to the flat dollar amount, no cents there are 100 cents in 1 dollar so rounding to the nearest cent, would be rounding to the nearest one hundredth

OpenStudy (anonymous):

419.20

OpenStudy (jdoe0001):

well 419.21538 ^ the 5 will "tip over" the 21 to 22

OpenStudy (anonymous):

so its 419.22

OpenStudy (jdoe0001):

yeap